OpenStudy (anonymous):

shown that the points A(-1,2),B(5,2) and C(2,5) are the vertices of an isosceles triangle find the area of ABC

OpenStudy (anonymous):

I found the answer to AB,BC,and AC now can anyone tell me how to find the area

OpenStudy (anonymous):

area = 1/2 * base*height

OpenStudy (anonymous):

The answer to AB=6,BC=\/18,AC=\/18

OpenStudy (anonymous):

It doesn't really matter what way you find the base or height

OpenStudy (anonymous):

I know the formula to find the area how am i suppose to find the hieght?

OpenStudy (anonymous):

find the midpoint of AB, call that D, then find the length of CD Area = CD x AB x 1/2

OpenStudy (anonymous):

the midpoint is (2,2)

OpenStudy (anonymous):

ok so I get the height 3

OpenStudy (anonymous):

What the hell just happened?? I Answered my own question I found the answer..I'm a genius..lol Thanx for ya'll help

OpenStudy (anonymous):

wait a sec

OpenStudy (apoorvk):

You found the 3 distances right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it says find the length perpendicular form a to bc

OpenStudy (anonymous):

from*

OpenStudy (apoorvk):

No you don't need that much trouble for the area. Ever heard of Heron's Formula?

OpenStudy (anonymous):

no...If its easier that way can you please explain it to me?

OpenStudy (apoorvk):

Ohkay, so if the lengths of the 3 sides of a triangle are a,b and c, then, semi-perimeter (half perimeter) = \(\large s = \frac{a+b+c}{2}\) Then, area 'A' of the triangle is: \[A = \sqrt{s(s-a)(s-b)(s-c)}\] (believe it's easier than it looks)

OpenStudy (apoorvk):

*believe me

OpenStudy (anonymous):

and one more thing sometimes they ask you to find the length perpendicular ..like from A to BC

OpenStudy (anonymous):

let me look at the formula for a moment

OpenStudy (apoorvk):

Ohkay if they have asked you to find the length of the perpendicular, then in this particular case, the perpendicular meets the opposite side at its mid-point (since this is an isosceles triangle)

OpenStudy (anonymous):

what does "large s" and "frac" I know but just making sure

OpenStudy (anonymous):

stand for*

OpenStudy (apoorvk):

Oh seems like the codes haven't processed, just refresh your tab once, that would make it alright.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok now their good I'm gonna look at em for a moment

OpenStudy (anonymous):

I've pretty much got it but what does the "s" stand for?

OpenStudy (apoorvk):

s = (a+b+c)/2 simply this^. (called the semi-perimeter)

OpenStudy (anonymous):

wow that pretty much solves the whole thing in a bit thanx again your incredible

OpenStudy (anonymous):

ok so lets look at the perpendicular question what will i have to do if its not an isosceles triangle ?

OpenStudy (anonymous):

Thanx for your Help I'm gonna go ahead and give my test bye

OpenStudy (apoorvk):

Hmm, that's difficult, let me think about ti and reply. All the best for your test man! :]

OpenStudy (anonymous):

Thank you